OpenStudy (anonymous):

the sum of 3 consecutive even integers is 156. find the smallest. a.52 b.46 c.50 d.48

OpenStudy (anonymous):

consecutive even intergers means thing like 2,4,6,8 etc.. so 3 consecutive intergers that equal 156 would be 50,52, and 54. the least being 50

OpenStudy (anonymous):

x+x+2+x+4=156 x=50

OpenStudy (lalaly):

since they are even numbers they should be 2, 4 , 6 ,... so there is a difference of two between two consecutive numbers so n+(n+2)+(n+4)=156 3n+6=156 3n=150 n=50

OpenStudy (anonymous):

thanks so much

OpenStudy (lalaly):

=) ur welcome:D

OpenStudy (anonymous):

dang lalaly thats a way better way to explain it XD

OpenStudy (lalaly):

lol urs was good :D

OpenStudy (anonymous):

Let x be the middle integer. Solve the following for x: (x - 2) + x + (x + 2) =156 3x = 156 x = 52 x-2 = 50, the smallest ---------------------------------- the equation for maya's bicycle speed is:\[\frac{6}{\text{maya}-8}\text{=}\frac{10.8}{\text{maya}}\]\[\text{maya}=18 \text{ km}/h \]